8 Wisconsin Snowmobile Corp. is considering a switch to level production. Cost efficiencies would occur under...

8

Wisconsin Snowmobile Corp. is considering a switch to level production. Cost efficiencies would occur under level production, and aftertax costs would decline by $45,600, but inventory would increase by $380,000. Wisconsin Snowmobile would have to finance the extra inventory at a cost of 13.0 percent.  

a-1. Determine the extra cost or savings of switching over to level production.
Answer #1

The company will spend to finance extra inventory =380000*13% = 49400

a-1 Extra costs =45600-49400= -3800

The company should not take the project as it would incur extra costs

But at r =45600/380000 = 12% the project can be taken.